# Engagement, User Satisfaction, and the Amplification of Divisive Content on Social Media

### v5

@@ Acknowledgements
Primary Category of Content (Our Study) Weight Percent
Entertainment
424.54 29.24
Other
103.86 7.15
Politics
493.07 33.96
News
268.71 18.51
Hobbies
148.81 10.25
Work
13.02 0.90
Total 1452 100
Table 9: At the end of the study, we ask participants about the content shown to them: “What were the tweets we showed you today predominantly about? Select a maximum of two.”
Primary Reason for Using Twitter (Our Study) Weight Percent
A way to stay informed
293.91 36.46
Entertainment
397.41 49.31
Keeping me connected to other people
48.16 5.98
It’s useful for my job or school
28.18 3.50
Lets me see different points of view
16.52 2.05
A way to express my opinions
21.82 2.71
@@
Table 10: We ask participants about the primary reason they use Twitter: “What would you say is the main reason you use Twitter?”
@@
In this subsection, we present demographic distributions for the 1030 participants in the 2020 ANES study (?) who reported using Twitter at least a few times a week (only those who reported using Twitter at least a few times a week were allowed to participate in our study). We weighed each participant using the relevant post-stratified weights provided by ANES. For certain demographic questions, participants were not given the option to opt out in our study, but they were in the ANES study. Thus, for those cases, we do not include the ANES participants who opted out of the question.
Race (ANES Study) Weight Percent
White
712.31 63.33
Black or African American
98.15 8.73
Hispanic
173.54 15.43
Asian or Native Hawaiian/other Pacific Islander
69.51 6.18
American Indian/Alaska Native or Other
21.92 1.95
Multiple races, non-Hispanic
49.33 4.39
Total 1125 100
Sex (ANES Study) Weight Percent
Male
628.24 55.85
Female
496.60 44.15
Total 1125 100
Political Party (ANES Study) Weight Percent
Democrat
468.69 41.64
Republican
235.41 20.91
Independent
393.78 34.99
Other
27.67 2.46
Total 1125 100
Education Levels (ANES Study) Weight Percent
Some high school
38.93 3.45
High school graduate
438.65 38.85
Associate degree
90.81 8.04
Bachelor’s degree
358.36 31.74
Master’s degree or above
187.77 16.63
Other
14.67 1.30
Prefer not to answer
0.00 0.00
Total 1129 100
Table 11: Distribution of educational attainment in the 2020 ANES Twitter population. For comparison purposes, we bin the ANES education categories, which are more fine-grained than ours, into our coarser categories.
Age Group (ANES Study) Weight Percent
18-24 years old
276.06 25.12
25-34 years old
215.85 19.64
35-44 years old
216.23 19.68
45-54 years old
156.59 14.25
55-64 years old
142.85 13.00
65-74 years old
66.98 6.10
75 years or older
24.41 2.22
Total 1099 100
Table 12: Distribution of ages in the 2020 ANES Twitter population. For comparison purposes, we group ages to match the age groups that we use in our study.
S2.3 Comparison to ANES 2020 demographic data
Finally, we present figures comparing our user demographics to Twitter users in the ANES 2020 who used Twitter at least a few times a week. The primary differences are that our population is younger (53 percent of our study are aged 18-34 years old, compared to 45 percent in the ANES study) and more likely to affiliate with the Democratic Party (56 percent Democrat in our study versus 41 percent in the ANES study).
Figure 4: Race of our study participants compared to Twitter users in the ANES 2020 study. We combine the race and ethnicity questions in our study into the summary categories used by ANES, shown in the figure.
Figure 5: Gender of our study participants compared to the sex of Twitter users in the ANES 2020 study. Note that we collected gender but not sex and ANES collected sex but not gender. As such, we can only compare Male/Female against Men/Women.
Figure 6: Self-reported political party affiliations of our study participants compared to that of Twitter users in the ANES 2020 study.
Figure 7: The highest education level achieved by our study participants compared to Twitter users in the ANES 2020 study. We bin the ANES education categories, which are more fine-grained than ours, into our coarser categories.
Figure 8: Ages of our study participants compared to Twitter users in the ANES 2020 study. We bin ages from the ANES study to match our categories, shown in the figure.
@@
Unsurprisingly, the number of likes and retweets that each tweet has is much higher in the engagement-based timeline. And as expected, the reader is less likely to follow the authors of tweets shown by the engagement-based algorithm compared to the chronological ranking. Interestingly, the average number of links is almost halved, which is consistent with the results of a prior study that used eight sock puppets to audit properties of Twitter’s algorithm (?).
Finally, the algorithm does not necessarily favor the most “popular” accounts: authors shown by the algorithm have a lower median number of followers and are less likely to be verified. It is important to clarify that our research was conducted before Twitter’s changed its verification policy on April 1, 2023. The new policy limits verification (indicated by a blue checkmark on Twitter profiles) to Twitter Blue subscribers who pay for the service. Our data was collected under the previous policy which required users to be “active”, “notable”, and “authentic” to get verified status. Consequently, the verified users in our data set tend to be official organizational accounts, celebrities, journalists, or other well-known individuals.
@@
Next, in Table 16, we show which accounts were most and least amplified by Twitter’s algorithm, where the amplification of an individual account is simply calculated as the number of tweets by that author in users’ engagement-based timelines minus the number of tweets by that author in users’ chronological timelines. Consistent with prior findings (?), the accounts that were most de-amplified were news outlets. This is likely because news outlets post more frequently than ordinary accounts. Though the chronological timeline is susceptible to this kind of spam, the engagement-based timeline is not (?).
Elon Musk also stands out as a significant outlier, receiving higher amplification than any other account by a wide margin. Notably, during our study period, Platformer alleged that Twitter deployed code to artificially boost the reach of Musk’s tweets (?). Though Musk denied such claims (?), our data seems to suggest that he did receive an abnormally high amount of amplification, relative to other accounts.
@@
@Reuters
532
@elonmusk
1387
@nytimes
529
@POTUS
295
@FatKidDeals
407
@RonFilipkowski
210
@Wario64
365
@JackPosobiec
171
@CNN
365
@fasc1nate
152
@NBA
280
@nytimes
135
@Forbes
280
@barstoolsports
134
@FoxNews
266
@DailyLoud
133
@AP
251
@hodgetwins
126
@washingtonpost
250
@CNN
119
@POTUS
206
@NBA
115
@IGN
204
@stillgray
107
@videogamedeals
185
@vidsthatgohard
103
@TMZ
179
@ClownWorld_
102
@elonmusk
151
@catturd2
97
@business
151
@BornAKang
96
@TheOnion
148
@krassenstein
95
@barstoolsports
128
@Wario64
93
@PlayStation
127
@HumansNoContext
90
@NFL
126
@Dexerto
84
@JackPosobiec
123
@CalltoActivism
84
@guardian
122
@mmpadellan
82
@Independent
119
@NFL
82
@ABC
114
@joncoopertweets
81
@nypost
108
@jordanbpeterson
79
@@
@elonmusk
1236.0
@Reuters
-518.0
@RonFilipkowski
144.0
@nytimes
-394.0
@fasc1nate
119.0
@Wario64
-272.0
@DailyLoud
114.0
@Forbes
-265.0
@hodgetwins
102.0
@CNN
-246.0
@vidsthatgohard
97.0
@AP
-214.0
@stillgray
89.0
@FoxNews
-198.0
@POTUS
89.0
@washingtonpost
-188.0
@BornAKang
87.0
@NBA
-165.0
@ClownWorld_
83.0
@videogamedeals
-156.0
@krassenstein
83.0
@IGN
-148.0
@HumansNoContext
80.0
@TMZ
-148.0
@Dexerto
73.0
@business
-139.0
@rawsalerts
68.0
@TheOnion
-136.0
@joncoopertweets
67.0
@guardian
-117.0
@buitengebieden
60.0
@Independent
-108.0
@Acyn
59.0
@PlayStation
-107.0
@mmpadellan
58.0
@TheEconomist
-85.0
@EndWokeness
58.0
@people
-82.0
@CollinRugg
58.0
@ABC
-79.0
@historyinmemes
57.0
@WSJ
-79.0
@CalltoActivism
53.0
@BBCWorld
-69.0
@libsoftiktok
52.0
@netflix
-68.0
@FightHaven
49.0
@HuffPost
-65.0
@JackPosobiec
48.0
@thehill
-63.0
@@
First, we show the distribution of survey responses for questions gauging the author and reader’s emotions (sad, happy, anxious, and angry), considering both the distribution for tweets overall and for only political tweets, in Figure 10 and Tables 17-20. For further ease of interpretation, in Figure 11, we also show the mean level of author and reader emotions in the chronological and engagement-based timelines after binarizing responses. In particular, if a reader responds “Not at all” to an emotion, that is coded as zero, while any response between “Slightly” and “Extremely” is coded as one. Notably, there was a large increase in tweets expressing anger between the engagement and chronological timeline. In particular, 62 percent of political tweets in the engagement timeline expressed anger, compared to 52 percent in the chronological timeline.
@@
Next, we show the distribution of survey responses for political questions (political leaning, out-group animosity, in-group perception, and out-group perception). We show the distributions for users overall (Figure 12, Tables 23-25) as well as separately for left-leaning (Figure 13, Tables 26-28) and right-leaning users (Figure 14, Tables 29-31).
@@
Far left (-2)
8.48
10.64
Left (-1)
34.20
32.70
Moderate (0)
41.27
33.94
Right (1)
12.20
15.28
Far right (2)
3.85
7.44
@@
Much worse (-2)
0.87
0.64
11.97
15.63
Worse (-1)
5.27
5.91
21.90
22.51
The same as before (0)
72.56
69.91
63.58
59.21
Better (1)
15.00
16.07
2.33
2.07
Much better (2)
6.30
7.47
0.22
0.58
@@
Far left (-2)
9.77
11.42
Left (-1)
39.09
39.05
Moderate (0)
42.76
33.85
Right (1)
6.12
9.54
Far right (2)
2.26
6.15
@@
Much worse (-2)
0.64
0.58
11.30
15.84
Worse (-1)
4.74
4.91
22.59
22.78
The same as before (0)
72.76
70.76
64.19
59.49
Better (1)
15.19
16.16
1.82
1.48
Much better (2)
6.68
7.59
0.10
0.41
@@
Far left (-2)
4.11
8.19
Left (-1)
17.37
12.65
Moderate (0)
36.96
34.74
Right (1)
32.46
33.17
Far right (2)
9.10
11.24
@@
Much worse (-2)
2.01
1.21
14.35
14.94
Worse (-1)
7.02
9.06
19.44
21.29
The same as before (0)
71.97
67.17
61.32
58.40
Better (1)
14.07
15.47
3.97
3.91
Much better (2)
4.92
7.09
0.93
1.46
@@
Here, we explore how the outcomes measured change based on the position of the tweet in each timeline (Figures 16-18). A priori, we expect that the outcomes in the chronological timeline have no correlation with tweet position since the chronological timeline is simply ordered by time. However, it would be reasonable to hypothesize that higher-ranked tweets in the engagement-based timeline would exhibit differences from lower-ranked tweets from the timeline. That being said, we did not find any statistically significant differences between the high-ranked tweets and low-ranked tweets, even in the engagement-based timeline. However, it is important to note that we only consider the top 10 tweets in the timeline. It could be the case that if we considered more tweets, we would find larger differences between high-ranked tweets and low-ranked tweets.
@@
In our main analysis, all outcomes are measured through survey questions given to readers; this includes both outcomes that explicitly ask about the reader’s perspective (i.e., the reader’s emotions, in-group perceptions, out-group perceptions, and stated preferences) as well as outcomes that pertain to the tweet itself (the author’s emotions, the ideological leaning of the tweet, and whether the tweet contains out-group animosity). Here, we elicit labels for the tweet-based outcomes from GPT-4 (?) and report effects based on these machine learning labels (Table 34 and Figure 22).
@@ 1. How is @{author_main} feeling in their tweet?
    - Angry
    - Anxious
    - Happy
    - Sad
@@
    - Angry
    - Anxious
    - Happy
    - Sad
@@ 6. How is @{author_other} feeling in their tweet?
    - Angry
    - Anxious
    - Happy
    - Sad
@@ 10. Is @{author_other}’s tweet expressing anger, frustration, or hostility
Ranking by stated preferences (SP) reduced the amount of partisan animosity, relative to the engagement-based timeline. However, this primarily occurred through a reduction in animosity towards the reader’s in-group, rather than animosity towards the reader’s out-group. In other words, there was an asymmetry in which readers tended to be tolerant of animosity towards their out-group but not their own in-group. This asymmetry caused a slight echo chamber effect, in which the SP timeline had the highest proportion of in-group content and the lowest proportion of out-group content (relative to both the engagement and chronological timeline). Is it possible to satisfy users’ stated preferences without amplifying echo chambers? To investigate this, we considered a variant of the SP timeline that used the presence of out-group animosity to break ties between tweets. We call this new timeline the SP-OA timeline.
@@
In conclusion, the SP-OA timeline has high satisfaction of users’ stated preferences, mitigates the amplification of divisive content, and does so without reinforcing echo chambers.

### v6

@@ Acknowledgements
Primary Category of Content (Our Study) $\mathbf{n}$ Percent
Entertainment 416 28.75
Other 108 7.46
Politics 483 33.38
News 272 18.80
Hobbies 152 10.50
Work 16 1.11
Total 1447 100
Table 8: At the end of the study, we ask participants about the content shown to them: ”What were the tweets we showed you today predominantly about? Select a maximum of two.”
Primary Reason for Using Twitter (Our Study) $\mathbf{n}$ Percent
A way to stay informed 293 36.35
Entertainment 393 48.76
Keeping me connected to other people 53 6.58
It’s useful for my job or school 28 3.47
Lets me see different points of view 18 2.23
A way to express my opinions 21 2.61
@@
Table 9: We ask participants about the primary reason they use Twitter: ”What would you say is the main reason you use Twitter?”
@@
In this subsection, we present demographic distributions for the 1030 participants in the 2020 ANES study [3] who reported using Twitter at least a few times a week (only those who reported using Twitter at least a few times a week were allowed to participate in our study). For certain demographic questions, participants were not given the option to opt out in our study, but they were in the ANES study. Thus, for those cases, we do not include the ANES participants who opted out of the question.
Comparing the demographics of users in the ANES 2020 study to those of users in our study, the largest differences are that our population is younger (53 percent of our study are aged 18-34 years old, compared to 33 percent in the ANES study) and more likely to affiliate with the Democratic Party (56 percent Democrat in our study versus 43 percent in the ANES study). That said, all distributions over demographic attributes are significantly different (when using a chi-squared test), apart from sex/gender.
Race/Ethnicity ANES Our Study
$n$ Percent $n$ Percent
White 729 71.19 537 66.63
Black or African American 71 6.93 89 11.04
Hispanic 113 11.04 81 10.05
Asian or Native Hawaiian/other Pacific Islander 49 4.79 63 7.82
American Indian/Alaska Native or Other 25 2.44 8 0.99
Multiple races, non-Hispanic 37 3.61 28 3.47
Total 1024 100.00 806 100.00
[…]
Sex/Gender ANES Our Study
$n$ Percent $n$ Percent
Male/Man 558 54.39 410 52.50
Female/Woman 468 45.61 371 47.50
Total 1026 100.00 781 100.00
[…]
Political Party ANES Our Study
$n$ Percent $n$ Percent
Democrat 439 42.75 451 55.96
Republican 222 21.62 111 13.77
Independent 338 32.91 205 25.43
Other 28 2.73 39 4.84
Total 1027 100.00 806 100.00
[…]
Education Level ANES Our Study
$n$ Percent $n$ Percent
Some high school 20 1.94 5 0.62
High school graduate 285 27.67 242 30.02
Associate degree 97 9.42 102 12.66
Bachelor’s degree 341 33.11 310 38.46
Master’s degree or above 275 26.70 128 15.88
Other 12 1.17 17 2.11
Prefer not to answer 0 0.00 2 0.25
Total 1030 100.00 806 100.00
[…]
Age Group ANES Our Study
$n$ Percent $n$ Percent
18-24 years old 131 13.14 120 14.89
25-34 years old 201 20.16 307 38.09
35-44 years old 235 23.57 211 26.18
45-54 years old 164 16.45 93 11.54
55-64 years old 137 13.74 51 6.33
65-74 years old 98 9.83 23 2.85
75 years or older 31 3.11 1 0.12
Total 997 100.00 806 100.00
[…]
@@
Unsurprisingly, the number of likes and retweets that each tweet has is much higher in the engagement-based timeline. And as expected, the reader is less likely to follow the authors of tweets shown by the engagement-based algorithm compared to the chronological ranking. Interestingly, the average number of links is almost halved, which is consistent with the results of a prior study that used eight sock puppets to audit properties of Twitter’s algorithm [6].
The algorithm does not necessarily favor the most “popular” accounts: authors shown by the algorithm have a lower median number of followers and are less likely to be verified. It is important to clarify that our research was conducted before Twitter’s changed its verification policy on April 1, 2023. The new policy limits verification (indicated by a blue checkmark on Twitter profiles) to Twitter Blue subscribers who pay for the service. Our data was collected under the previous policy which required users to be “active”, “notable”, and “authentic” to get verified status. Consequently, the verified users in our data set tend to be official organizational accounts, celebrities, journalists, or other well-known individuals.
Finally, Figure 5 shows the distribution of the number of political tweets in the user timelines. There was little difference between the distribution of political tweets in the engagement-based timelines and the chronological timelines. About 70% of participants’ timelines contained at least one political tweet (out of approximately ten tweets total).
@@
Figure 5: The distribution of political tweets in user timelines. The number of political tweets in each timeline is calculated using the participant’s response to the binary question, “Is [@author-handle]’s tweet about a political or social issue?”. Notably, about 30% of participants’ timelines contain no political tweets.
@@
Next, in Table 18, we show which accounts were most and least amplified by Twitter’s algorithm, where the amplification of an individual account is simply calculated as the number of tweets by that author in users’ engagement-based timelines minus the number of tweets by that author in users’ chronological timelines. Consistent with prior findings [6], the accounts that were most de-amplified were news outlets. This is likely because news outlets post more frequently than ordinary accounts. Though the chronological timeline is susceptible to this kind of spam, the engagement-based timeline is not [40]. Thus, whether users are using their chronlogical or engagement-based timeline may have a large influence on their exposure to traditional news outlets [53].
Elon Musk also stands out as a significant outlier, receiving higher amplification than any other account by a wide margin. Notably, during our study period, Platformer alleged that Twitter deployed code to artificially boost the reach of Musk’s tweets [45]. Though Musk denied such claims [23], our data seems to suggest that he did receive an abnormally high amount of amplification, relative to other accounts.
@@
@Reuters 532 @elonmusk 1387
@nytimes 529 @POTUS 295
@FatKidDeals 407 @RonFilipkowski 210
@Wario64 365 @JackPosobiec 171
@CNN 365 @fasc1nate 152
@NBA 280 @nytimes 135
@Forbes 280 @barstoolsports 134
@FoxNews 266 @DailyLoud 133
@AP 251 @hodgetwins 126
@washingtonpost 250 @CNN 119
@POTUS 206 @NBA 115
@IGN 204 @stillgray 107
@videogamedeals 185 @vidsthatgohard 103
@TMZ 179 @ClownWorld_ 102
@elonmusk 151 @catturd2 97
@business 151 @BornAKang 96
@TheOnion 148 @krassenstein 95
@barstoolsports 128 @Wario64 93
@PlayStation 127 @HumansNoContext 90
@NFL 126 @Dexerto 84
@JackPosobiec 123 @CalltoActivism 84
@guardian 122 @mmpadellan 82
@Independent 119 @NFL 82
@ABC 114 @joncoopertweets 81
@nypost 108 @jordanbpeterson 79
@@
@elonmusk 1236.0 @Reuters -518.0
@RonFilipkowski 144.0 @nytimes -394.0
@fasc1nate 119.0 @Wario64 -272.0
@DailyLoud 114.0 @Forbes -265.0
@hodgetwins 102.0 @CNN -246.0
@vidsthatgohard 97.0 @AP -214.0
@stillgray 89.0 @FoxNews -198.0
@POTUS 89.0 @washingtonpost -188.0
@BornAKang 87.0 @NBA -165.0
@ClownWorld_ 83.0 @videogamedeals -156.0
@krassenstein 83.0 @IGN -148.0
@HumansNoContext 80.0 @TMZ -148.0
@Dexerto 73.0 @business -139.0
@rawsalerts 68.0 @TheOnion -136.0
@joncoopertweets 67.0 @guardian -117.0
@buitengebieden 60.0 @Independent -108.0
@Acyn 59.0 @PlayStation -107.0
@mmpadellan 58.0 @TheEconomist -85.0
@EndWokeness 58.0 @people -82.0
@CollinRugg 58.0 @ABC -79.0
@historyinmemes 57.0 @WSJ -79.0
@CalltoActivism 53.0 @BBCWorld -69.0
@libsoftiktok 52.0 @netflix -68.0
@FightHaven 49.0 @HuffPost -65.0
@JackPosobiec 48.0 @thehill -63.0
@@
First, we show the distribution of survey responses for questions gauging the author and reader’s emotions (sad, happy, anxious, and angry), considering both the distribution for tweets overall and for only political tweets, in Figure 6 and Tables 19-22. For further ease of interpretation, in Figure 7, we also show the mean level of author and reader emotions in the chronological and engagement-based timelines after binarizing responses. In particular, if a reader responds “Not at all” to an emotion, that is coded as zero, while any response between “Slightly” and “Extremely” is coded as one. Notably, there was a large increase in tweets expressing anger between the engagement and chronological timeline. In particular, 62 percent of political tweets in the engagement timeline expressed anger, compared to 52 percent in the chronological timeline.
@@
Next, we show the distribution of survey responses for political questions (political leaning, out-group animosity, in-group perception, and out-group perception). We show the distributions for users overall (Figure 8, Tables 25-27) as well as separately for left-leaning (Figure 9, Tables 28-30) and right-leaning users (Figure 10, Tables 31-33).
@@
Far left (-2) 8.48 10.64
Left (-1) 34.20 32.70
Moderate (0) 41.27 33.94
Right (1) 12.20 15.28
Far right (2) 3.85 7.44
@@
Much worse (-2) 0.87 0.64 11.97 15.63
Worse (-1) 5.27 5.91 21.90 22.51
The same as before (0) 72.56 69.91 63.58 59.21
Better (1) 15.00 16.07 2.33 2.07
Much better (2) 6.30 7.47 0.22 0.58
@@
Far left (-2) 9.77 11.42
Left (-1) 39.09 39.05
Moderate (0) 42.76 33.85
Right (1) 6.12 9.54
Far right (2) 2.26 6.15
@@
Much worse (-2) 0.64 0.58 11.30 15.84
Worse (-1) 4.74 4.91 22.59 22.78
The same as before (0) 72.76 70.76 64.19 59.49
Better (1) 15.19 16.16 1.82 1.48
Much better (2) 6.68 7.59 0.10 0.41
@@
Far left (-2) 4.11 8.19
Left (-1) 17.37 12.65
Moderate (0) 36.96 34.74
Right (1) 32.46 33.17
Far right (2) 9.10 11.24
@@
Much worse (-2) 2.01 1.21 14.35 14.94
Worse (-1) 7.02 9.06 19.44 21.29
The same as before (0) 71.97 67.17 61.32 58.40
Better (1) 14.07 15.47 3.97 3.91
Much better (2) 4.92 7.09 0.93 1.46
@@
Here, we explore how the outcomes measured change based on the position of the tweet in each timeline (Figures 12-14). A priori, we expect that the outcomes in the chronological timeline have no correlation with tweet position since the chronological timeline is simply ordered by time. However, it would be reasonable to hypothesize that higher-ranked tweets in the engagement-based timeline would exhibit differences from lower-ranked tweets from the timeline. That being said, we did not find any statistically significant differences between the high-ranked tweets and low-ranked tweets, even in the engagement-based timeline. However, it is important to note that we only consider the top 10 tweets in the timeline. It could be the case that if we considered more tweets, we would find larger differences between high-ranked tweets and low-ranked tweets.
@@
In our main analysis, all outcomes are measured through survey questions given to readers; this includes both outcomes that explicitly ask about the reader’s perspective (i.e., the reader’s emotions, in-group perceptions, out-group perceptions, and stated preferences) as well as outcomes that pertain to the tweet itself (the author’s emotions, the ideological leaning of the tweet, and whether the tweet contains out-group animosity). Here, we elicit labels for the tweet-based outcomes from GPT-4 [39] and report effects based on these machine learning labels (Table 36 and Figure 18).
@@ 1. How is @{author_main} feeling in their tweet?
- Angry
- Anxious
- Happy
- Sad
@@
- Angry
- Anxious
- Happy
- Sad
@@ 6. How is @{author_other} feeling in their tweet?
- Angry
- Anxious
- Happy
- Sad
@@ 10. Is @{author_other}’s tweet expressing anger, frustration, or hostility
Ranking by stated preferences (SP) reduced the amount of partisan animosity, relative to the engagement-based timeline. However, this primarily occurred through a reduction in animosity towards the reader’s in-group, rather than animosity towards the reader’s out-group. In other words, there was an asymmetry in which readers tended to be tolerant of animosity towards their out-group but not their own in-group. The SP timeline had the highest proportion of in-group content and the lowest proportion of out-group content (relative to both the engagement and chronological timeline). Is it possible to satisfy users’ stated preferences without inducing in-group bias? To investigate this, we considered a variant of the SP timeline that used the presence of out-group animosity to break ties between tweets. We call this new timeline the SP-OA timeline.
@@
In conclusion, the SP-OA timeline has high satisfaction of users’ stated preferences, mitigates the amplification of divisive content, and does so without reinforcing in-group bias.